# Sink-free orientations: a local sampler with applications

### v1

@@ Abstract.
For sink-free orientations in graphs of minimum degree at least $3$, we show that there is a deterministic approximate counting algorithm that runs in time $O((n^{73}/\varepsilon^{72})\log(n/\varepsilon))$, a near-linear time sampling algorithm, and a randomised approximate counting algorithm that runs in time $O((n/\varepsilon)^{2}\log(n/\varepsilon))$, where $n$ denotes the number of vertices of the input graph and $10<\varepsilon<$ 1 is the desired accuracy. All three algorithms are based on a local implementation of the sink popping method (Cohn, Pemantle, Propp, 2002) under the partial rejection sampling framework (Guo, Jerrum, Liu, 2019).
@@ 1.1. Our contribution and technique overview
One much less obvious alternative approach to FPTAS is via the connection of the local lemma. In particular, because SFOs form extremal instances, their number can be computed via the independence polynomial evaluated at negative weights on the dependency graph. (We also see this fact in Section 4.1.) Normally this approach would not be efficient, because the dependency graph is usually exponentially large (for example for all-terminal reliability), but in the case of SFOs, the dependency graph is just the input graph itself. There are more than one FPTASes [35, 24] for the independence polynomial at negative weights. However, neither appears able to recover Theorem 1.1. We provide a detailed discussion in Section 4.2. Here, let us briefly mention that the algorithm by Patel and Regts [35] requires a uniform bound on the weights, which cannot handle the mixture of high and low degree vertices. On the other hand, with the minimum degree $\geq 3$ assumption, the probability vector for SFOs is within the so-called Shearer’s region [36]. This makes the algorithm by Harvey, Srivastava, and Vondrák [24] more relevant. However, the running time of their algorithm has the form $(n/\varepsilon)^{O(\log d)}$, [^3] To be more precise, the hidden constant in the exponent is linear in the inverse of the multiplicative “slack” of how close the evaluated point is to the boundary of Shearer’s region. For SFOs, when constant degree vertices are present, the slack is a constant. where $d$ is the maximum degree of the graph. Thus, in the setting of Theorem 1.1, their algorithm runs in quasi-polynomial time instead.
@@ 4.2. Independence polynomial at negative weights
There are more than one FPTASes that can efficiently approximate the independence polynomial at negative weights. The algorithm by Patel and Regts [35] requires that $\left|x_{u}\right|<\frac{(d-1)^{d-1}}{d^{d}}$ for all $u\in V$, where $d$ is the maximum degree. Thus, it does not apply to the case of SFOs when there are vertices of degree, say, both 3 and 4. As the threshold is decreasing in $d$, it actually does not apply to any graph with constant degree vertices together with vertices of unbounded degrees.
The other algorithm, by Harvey, Srivastava, and Vondrák [24], on the contrary, can work within Shearer’s region, and is thus more relevant. To explain Shearer’s region, let us abuse the notation slightly and extend the definition in (4) to a function $q_{J}(\mathbf{x})=\sum\limits_{I\in\textnormal{{Ind}}(G),I\subseteq J}\prod\limits_{u\in I}x_{u}$ to take an input weight vector $\mathbf{x}$. Then, a vector $\mathbf{p}$ is in Shearer’s region if and only if $q_{S}(-\mathbf{p})>0$ for all $S\subseteq V$. Lemma 4.1 implies that the probability vector for SFOs is in Shearer’s region. Moreover, we say a vector $\mathbf{p}$ has slack $\alpha$ if $(1+\alpha)\mathbf{p}$ is in Shearer’s region. For a vector $\mathbf{x}$ with slack $\alpha$, the algorithm in [24] approximates $q_{G}(\mathbf{x})$ in time $(n/(\alpha\varepsilon))^{O(\log d/\sqrt{\alpha})}$. It also does not recover Theorem 1.1 as the slack is a constant when constant degree vertices exist. If, in the meantime, some other vertices have unbounded degrees, the algorithm in [24] runs in quasi-polynomial time.

### v2

@@ Abstract.
For sink-free orientations in graphs of minimum degree at least $3$, we show that there is a deterministic approximate counting algorithm that runs in time $O((n^{73}/\varepsilon^{72})\log(n/\varepsilon))$, a near-linear time sampling algorithm, and a randomised approximate counting algorithm that runs in time $O((n/\varepsilon)^{2}\log(n/\varepsilon))$, where $n$ denotes the number of vertices of the input graph and $10<\varepsilon<$ 1 is the desired accuracy. All three algorithms are based on a local implementation of the sink popping method (Cohn, Pemantle, and Propp, 2002) under the partial rejection sampling framework (Guo, Jerrum, and Liu, 2019).
@@ 1.1. Our contribution and technique overview
One much less obvious alternative approach to FPTAS is via the connection of the local lemma. In particular, because SFOs form extremal instances, their number can be computed via the independence polynomial evaluated at negative weights on the dependency graph. (We also see this fact in Section 4.1.) Normally this approach would not be efficient, because the dependency graph is usually exponentially large (for example for all-terminal reliability), but in the case of SFOs, the dependency graph is just the input graph itself. There are more than one FPTASes [35, 24] for the independence polynomial at negative weights. However, neither appears able to recover Theorem 1.1. With the minimum degree $\geq 3$ assumption, the probability vector for SFOs is within the so-called Shearer’s region, where both algorithms apply. [^3] In [35], only a uniform bound is stated, but one can introduce a scaling variable $t$ and make a new polynomial in $t$, so that their algorithm works in the Shearer’s region. The downside is that the running time of both algorithms has the form $(n/\varepsilon)^{O(\log d)}$, [^4] To be more precise, the hidden constants in the exponents decrease in the multiplicative “slack” of how close the evaluated point is to the boundary of Shearer’s region. For SFOs, when constant degree vertices are present, the slack is a constant, and so are the hidden constants in the exponents. where $d$ is the maximum degree of the graph. Thus, in the setting of Theorem 1.1, these algorithms run in quasi-polynomial time instead. A more detailed discussion is given in Section 4.2.
@@ 4.2. Independence polynomial at negative weights
There are more than one FPTASes [35, 24] that can efficiently approximate the independence polynomial at negative weights. These algorithms work in the so-called Shearer’s region [36]. To explain Shearer’s region, let us abuse the notation slightly and extend the definition in (4) to a function $q_{J}(\mathbf{x})=\sum\limits_{I\in\textnormal{{Ind}}(G),I\subseteq J}\prod\limits_{u\in I}x_{u}$ to take an input weight vector $\mathbf{x}$. Then, a vector $\mathbf{p}$ is in Shearer’s region if and only if $q_{S}(-\mathbf{p})>0$ for all $S\subseteq V$. Lemma 4.1 implies that the probability vector for SFOs is in Shearer’s region. Moreover, we say a vector $\mathbf{p}$ has slack $\alpha$ if $(1+\alpha)\mathbf{p}$ is in Shearer’s region. For a vector $\mathbf{x}$ with slack $\alpha$, the algorithm by Patel and Regts [35] $\varepsilon$-approximates $q_{G}(\mathbf{x})$ in time $(n/\varepsilon)^{O(\log d/\alpha)}$, and the algorithm by Harvey, Srivastava, and Vondrák [24] runs in time $(n/(\alpha\varepsilon))^{O(\log d/\sqrt{\alpha})}$, where $d$ is the maximum degree of the graph. They do not recover Theorem 1.1 as the slack is a constant when constant degree vertices exist. If, in the meantime, some other vertices have unbounded degrees, these algorithms run in quasi-polynomial time instead.
@@ 5. Concluding remarks
Acknowledgement
We thank Guus Regts for helpful comments on an earlier version of this paper.